The condensed financial statements of Jenner Corporation for 2011 are presented below.

Malli Burton Malli Burton

Balance Sheet Income Statement

December 31, 2014 For the Year Ended December 31, 2014

Assets Revenues $500,000

Current assets Expenses

Cash and short-term Cost of goods sold 255,000

investments $ 15,000 Selling and administrative

Accounts receivable 17,500 expenses 170,000

Inventories 35,000 Interest expense 12,500

Total current assets 67,500 Total expenses 437,500

Property, plant, and Income before income taxes 62,500

equipment (net) 182,500 Income tax expense 25,000

Total assets $250,000 Net income $ 37,500

Liabilities and Stockholders' Equity

Current liabilities $ 25,000

Long-term liabilities 95,000

Stockholders' equity 130,000

Total liabilities and

stockholders' equity $250,000

Additional data as of December 31, 2013: Inventory = $31,000; Total assets = $210,000; Stockholders' equity = $140,000.

Instructions: Compute the following ratios for 2014 showing supporting calculations.

(a) Current ratio = .

(b) Debt to total assets ratio = .

(c) Times interest earned = .

(d) Inventory turnover = .

(e) Profit margin = .

(f) Return on stockholders' equity = .

(g) Return on assets = .
Answer #1
Jenner Corporation
a) Current Ratio
Current Assets/Current Liabilites
Current Assets
Cash & Short term Investment $                  15,000.00
Accounts Receivable $                  17,500.00
Inventories $                  35,000.00
Total Current Assets=(A) $                  67,500.00
Current Liabities
Total Current Liabilites=(B) $                  25,000.00
Current Ratio=(A)/(B) 2.70:1
b) Debt to Assets Ratio=Total Debt/Total Assets
Total Debt=Current liabilites+ Long Term liabilities=($25000+$95000)=(A) $              1,20,000.00
Total Assets=(B) $              2,50,000.00
Debt Ratio=Total Debt/Total Assets=(A)/(B)=($16238/$28362) 0.48
c ) Times Interest Earned=Income before interest and inome tax/Interest
Income before interest and income tax=($62500+$12500)=(A) $                  75,000.00
Interest=(B) $                  12,500.00
Times Interest Earned=(A)/(B) 6.00 Times
d) Inventory Turnover
Cost of goods sold=(A) $              2,55,000.00
Beginning Inventoery $                  31,000.00
Ending Inventory $                  35,000.00
Average Inventory=(Beginning Inventory+Ending Inventory/2)=($31000+$35000)/2=(B) $                  33,000.00
Accounts Receivable Turnover=Sales/Average accounts receivable=(A)/(B) 7.73 Times
e) Profit Margin=Net Income after tax/Sales
Net Income after tax=(A) $                  37,500.00
Sales=(B) $              5,00,000.00
Profit Margin=Net Income after tax/Sales=(A)/(B) 7.50%
f) Return on Stockholder's Equity
Net Income after tax/Total Stockholders equity
Net Income after tax=(A) $                  37,500.00
Total Stockholders Equity=(B) $              1,30,000.00
Return on Stockholder's Equity=(A)/(B) 28.85%
OR
Return on Average stockholder's Equity
Net Income after tax/Average Stockholders equity
Net Income after tax=(A) $                  37,500.00
Beginning Equity $              1,40,000.00
Ending Equity $              1,30,000.00
Average Stockholder's equity=(Beginning Equity+Ending Eqity)/2=($140000+$130000)/2=(B) $              1,35,000.00
Return on Average stockholder's Equity=(A)/(B) 27.78%
g) Return on Assets
Return on total assets=Earnings before interest and income tax/Total Assets
Earnings before interest and income tax($62500+$12500)=(A) $                  75,000.00
Total Assets=(B) $              2,50,000.00
Return on Total Assets=(A)/(B) $                             0.30
OR
Return on Total Assets
Net Income before interest and tax/Average Total Assets
NEt Income before interest and tax=(A) $                  75,000.00
Beginning Total Assets $              2,10,000.00
Ending Total Assets $              2,50,000.00
Average Total Assets=(Beginning Total Assets+Ending Total Assets)/2=($210000+$250000)/2=(B) $              2,30,000.00
Return on Total Assets=(A)/(B) 32.61%